A certain virus infects one in every 250 people. A test used to detect the virus in a person is positive 80% of the time when the person has the virus and 15% of the time when the person does not have the virus. (This 15% result is called a false positive)

Let A be the event "the person is infected" and B be the event "the person tests positive."

(a) Using Bayes' Theorem, when a person tests positive, determine the probability that the person is infected.

(b) Using Bayes' Theorem, when a person tests negative, determine the probability that the person is not infected.
Answer #1

ANSWER:

A certain virus infects one in every 250 people. A test used to detect the virus in a person is positive 80% of the time when the person has the virus and 15% of the time when the person does not have the virus. (This 15% result is called a false positive) Let A be the event "the person is infected" and B be the event "the person tests positive." (a) Using Bayes' Theorem, when a person tests positive, determine the probability that the person is infected. (b) Using Bayes' Theorem, when a person tests negative, determine the probability that the person is not infected.

Let a be the event the process is an injected then p(A) = 1/250  

A^{C} Be the event the person is not injected. Then p(A^{C}) =249/250 i.e (p(A^{C})= 1-p(A)

Let B be the event of test positive then p(B/A) = prob [that test is positive when the person is injected)

= 80%=0.8 and

p(B/A^{C}) = Prob(That test is positive when the person is injected )

=15% = 0.15

sol (A) : p(A/B) = prob(that person is infected when test is +ve)

= p(A).p(B/A) / (p(A).p(B/A) + p(A^{C})(B/A^{C})

= 0.004 * 0.8 / ((0.004*0.8) + (249/250)*0.15)

= 0.02096

sol b: where A^{C} be the event . The person is not injected B^{C_{}} be the event test is negative.

NOW p(B^{C_{}}/A) = 1- p(B/A) = 1- 80% = 1- 0.8 = 0.2 and

P(B^{C_{}}/ A^{C})= 1- P(B^{C_{}}/A) = 1 - 0.15 = 0.25

p(B^{C_{}}/ A^{C} ) =  p(A^{C}).p(B^{C_{}}/ A^{C} ) / (p(A^{C}).p(B^{C_{}}/ A^{C} ) + p(B^{C_{}})(/A)

= (249/250 ) * 0.25 / ((249/250)*0.25)+ (1/250)*0.2

= 0.99679